I have changed the MC2_YAW DC bias because the PZT1_YAW was railing.
I also realigned the steering mirrors in zig-zag path since the mode cleaner tended to resonate with higher order modes after I have changed the MC2 bias.
C1:SUS-MC2_YAW_COMM = -1.1548 => -1.1208
I found that the slow machine c1auxey, which controls and monitoring the ETMY suspension things, were not responding.
The machine responded to ping but I wasn't able to telnet to it.
I went down there and power-cycled it by keying the power of the VME rack, and then it came back and seems working properly.
I have no idea why it ran into such condition.
For some reason C1PEM doesn't seem to work anymore after a recompilation. It did recompile fine. We just changed some channel / subsystem names.
Tried reverting to the svn version. Doesn't work. Reboot C1SUS also no good.
It is fine again. Thanks Jamie.
Do we care about the AC? I thought what we care is the DC.
Right now the fluctuation of the green beat-note seems mostly covered by unknown noise which is relatively white.
The 2nd trial of the Y arm ALS noise budgeting :
(Removal of broad band noise)
I will add the dark noise of the broad-band beat-note PD and the MFD read out noise on the budget.
To reiterate: We changed the OSEM loop shape for MC1-MC3. Below in black is the old loop shape, which simulated pendulum response in there. In red is the new loop shape.
The differences are due to extra filter in C1:SUS-MC?_SUSPOS module 6,7,9
6: Elliptical LP @ 2.5Hz
7: Inverse Chebychev HP @0.3Hz
8: 1st order LP @ 10Hz
This is a trend for a day long showing the REFL11/55 demod signals, REFLDC (corresponding to the MC transmitted power) and the PSL booth temperatire.
There are sudden jumps in the REFL55_I and REFL11_Q signals around 5:00 AM this morning, also at the same time the temperature suddenly went up.
But the quality of the signal turned out to be not so good because the fluctuation is still within 1 bit of the ADCs,
we have to try it again with a bigger gain in the analog whitening circuit.
An RFAM measurement is ongoing
DO NOT CHANGE THE IFO ALIGNMENT UNTIL TOMORROW MORNING OR FURTHER NOTICE.
We're playing with the MC OAF, so we're actuating on MC2. Again, FYI.
I don't think I touched/adjusted/whatever anything, but I did open the PSL table ~5-10min ago to measure the size of the Kiwamu-Box, so if the RFAM stuff looks funny for a few minutes, it was probably me. Just FYI.
restarted Apache on Nodus for the SVN as per wiki instructions
Plus, MC has to be kept locked with the WFS.
Recovery from the power shutdown
- Turned on the raid disk of linux1.
- Woke linux1 up. No fsck this time.
- Woke up all the lab machines.
- Turned on all the electronics racks' AC powers
- Woke up fb and then front end machine (the raid for fb had been already up as I turned on the AC powers)
- Turned on all the electronics racks' DC powers (Sorensens, Kepcos, and etc.)
- Turned on the Marcnois which is driving the RF generation box.
- Woke up all the lasers (PSL and End lasers)
- Some burtrestoring (c1ioo, c1sus, c1susaux, c1msc, c1psl, c1iool0, c1auxey, c1auxex, c1oaf, c1pem)
- Ran autolockMC scripts on op340m => After relocking of PMC a lock of MC was acquired immediately.
- Turned on the PZT HV drivers.
- One of the Sorensens in 1X8 rack is showing the current limit sign. This is exactly the same situation as we saw before (#5592).
Currently it's off. It needs an investigation to find who is drawing such a large amount of current.
- C1SCX is not properly running. Rebooting the machine didn't help. This needs to be fixed.
The symptom is : (1) all the values are frozen in the screens. (2) the c1scx status screens shows NO SYNC sign. (3) however the timing board looks blinking happily.
- One of the VME rack on 1X3 is not showing the +/-15V green LED lights.
This is the one on very upper side of the rack, which contains the old c1lsc machine and c1iscaux2. If we are still using c1iscaux2, it needs to be fixed.
State condition: Vac Normal, CC1 ~8e-6 Torr
The IFO pressure peaked at 8.3 mTorr after 2days of not pumping.
We looked some more into the the OSEM signals and their coherence to the seismometer signals.
We were able to verify that the coherence OSEM sensor <-> seismometer signal goes down with increasing the OSEM gain. This seems to indicate that the OSEM FB add noise to the distance mirror <-> frame. We injected some noise into the OSEMs to see how the coherence behaves.
MC2 SUSPOS, 0.1Hz - 0.8Hz, 3mins each
Inj. amplitude Time(UTC) Note
- 21:35 Free swinging
- 21:42 Big LF OSEM gain
- 21:48 Small LF OSEM gain
150 21:56 -"-
300 22:00 -"-
900 22:05 -"-
High OSEM gain:
Low OSEM gain:
We left the filters that lower the OSEM gain below 0.3Hz on.
Could use some more detail on how this measurement was done. It looks like you used the SUSPOS signal with the mirror moving, however, this is not what we want. Of course, the SUSPOS with the mirror moving will always show the mirror motion because the OSEMs are motion sensors.
Instead, what we want is to project how the actual OSEM noise in the presence of no signal shows up as MC length. For that we should use the old traces of the OSEM noise with no magnets and then inject that spectrum of noise into the SUSPOS filter bank with all the loops running. We can then use this TF to estimate the projection of OSEM noise into the MC length.
As far as improving the damping filter, the 2.5 LP is not so hot since it doesn't help at low frequencies. Instead, one can compute the optimal filter for the SUSPOS feedback given the correct cost function. To first order this turns out to be the usual velocity damping filter but with a resonant gain at the pendulum resonance. This allows us to maintain the same gain at the pendulum mode but ~3x lower gain at other frequencies.
In the past, we had some issues with this due to finite cross-coupling with the angular loops. It would be interesting to see if we can use the optimal damping feedback now that the SUS DOFs have been diagonalized with the new procedure.
The measurement was done with the MC in lock and the OSEMS active.
1. I injected noise into MC1-3 SUSPOS_EXC at a level that domiated the SUSPOS output.
2. Then I calculated the coupling coefficients of the SUSPOS outputs to MC_F during the time the noise is injected.
3. Without noise injection I projected the SUSPOS outputs to MC_F by multiplying the coupling coefficients with the SUSPOS outputs.
All on 11-11-18. White noise inj. from 0.1Hz to 20Hz. Duration 4mins each.
DOF Amplitude(counts) Time(UTC)
MC1 200 22:08
MC2 25 22:25
MC3 25 22:50
As you say this does not show the dark / bright noise of the OSEMs. It shows the influence of the OSEMS output onto MC_F in normal operation of the MC. I would have expected that to be very low everywhere except at the pendulum resonance. Reason for that not to be true could either be the OSEMs having considerable gain off of the resonance, or noise intrinsic to the OSEMs knocking the mirrors around. Since we know the OSEM signal to MC_F TF we only need to compare the OSEM signal to OSEM noise to see the noise contribution to MC_F. We know from http://nodus.ligo.caltech.edu:8080/40m/5547 that the OSEM sensor bright noise is considerably below the OSEM signal above 0.1Hz in actual operation. We checked that the MC OSEM signals are above the noise in the reference above 0.1Hz by a factor 3-10.
We actually measured the cost function with the noise projection (valid to 10Hz). It's just the coupling coefficient, right?
I'm quite sure that this is not good: since MC2 can produce a signal in WFS1 and WFS2, it cannot be removed in this way from the actuation without introducing a significant cross-coupling between the MC_TRANS and WFS loops.
Really need loop TFs measured and compared with the model.
The WFS noise model will also show that we need to have a much lower UGF in the MCT loop since that sensor is just a DC QPD: it can never have as good of a sensing noise as a good WFS. In the current case with no Whitening, this is even more so.
Without adding significant amounts of noise to other WFS loops I have engaged the MC2_TRANS_PIT and YAW loops.
After several attempts to measure the system response and computing the output matrix, none of which gave any useful results, I gave up on that and decided to find three orthogonal actuation vectors which enable us to close the loops. So using the last good output matrix (below left side) as a template, I rounded it off to the nearest set of orthogonal vectors and arrived at the following matrix (right side):
I also decided that WFS1 and 2 need not drive MC2. This is just to decouple the loops and minimise cross-talk. This (albeit heuristic) matrix seems to work pretty well and the real matrix is probably quite close to it.
I show below the suppressed error signals after tweaking the gains a bit. The blue line is with no WFS, the green one with only WFS1 and 2 loops on, while the red is with all loops turned on. The WFS1Yaw and MC2_Trans_pit loops might benefit from a more careful study to determine a better output matrix.
That's right. The easier problem arises if we consider one of MC mirrors. The coherence between OSEM sensors and GUR1_X in free moving regime is equal to 0.9 at frequencies 0.1 - 1 Hz. But with local dumping coherence is 0.6. We have
Mirror -> Sensor -> Satellite Module -> Whitening -> ADC -> Computer -> DAC -> Dewhitening -> Satellite Module -> Actuator -> Mirror
Somewhere we produce noise that kills part of coherence. We can use this method with the injection of spectrum of noise into the SUSPOS filter bank only for one mirror and see how the coherence between OSEM sensor and GUR1_X will change. If the change is small, we deal with something else. It the coherence will change from 0.6 to ~0.4, than we have big OSEM noise.
It might be also the problem that the amplitude of COIL_OUT signal is ~25. If it is in counts we may have noise from DAC.
I installed the newly modified PDH box #17 and locked the Y-Arm.
I wasn't able to bring the REFL level down to the 30% that Kiwamu claimed to get, despite readjusting the alignment---I got ~40-45%. I attained a UGF of ~8 kHz, lower than the 20 kHz that Kiwamu said he got with the temporary setup, probably because the PDH box just isn't as fast. Despite that, it looks like the error suppression is actually better than before...
Here is an error spectrum:
I have to admit that this calibration is worthy of suspicion and should be done more rigorously. I simply used the measured UGF frequency and known servo TF and PZT actuator gain to estimate the optical response. I am pretty confident that it's accurate to within a factor of 3 or so.
I just figured out why the Newport 6000 isn't stabilizing the temperature. It is designed to drive a TEC, so that when the temperature is too high, it just applies a negative current. Of course, this doesn't work with a resistive heater; it just keeps heating it up more.
I'm not sure if Frank has actually used this with a restive heater before, but it doesn't appear that you can limit the low-current level or add an offset.
I used aluminum tape to attach the sensor and heater to the 40m's EOM, and we plugged in the controller. It seems to be kind of working. Zach figured out the GPIB output stuff, so we can talk to it remotely.
I stole the Prologix wireless GPIB interface from the SR785 that's down the Y-Arm temporarily. The address is 192.168.113.108. (Incidentally, I think some network settings have been changed since the GPIB stuff was initially configured. All the Prologix boxes have 131.215.X.X written on them, while they are only accessible via the 192.168.X.X addresses. Also, the 40MARS wireless router is only accessible from Martian computers at 192.168.113.226---not 220.127.116.11).
In any case, the Newport 6000 is controllable via telnet. I went through the remote RTD calibration process in the manual, by measuring the exact RTD resistance with an ohmmeter and entering it in. Despite this, when the TEC output is turned on, the heating way overshoots the entered set temperature. This is probably because the controller parameters (gain, etc.) are not set right. We have left it off for the moment.
Here are a couple command examples:
1. Turning on the TEC output
Some more info on this:
At these frequencies the pendulum should be quieter than the stacks. By quite a bit actually since there is the stack resonance at a couple Hz. 'Glueing' them together via the local control is not wise. We put an elliptic LP ( 2.5Hz, 4th order 6dB) into the C1:SUS-MC?_SUSPOS pathes and MC-F got better above 1Hz
Added an extra LP @ 10 Hz afterwards. Doesn't make a visible difference.
Now here is more stuff to consider.
1. The OSEMs glue the MC mirrors to the stacks
2. The pendulum TF should be 1
3. It shouldn't matter if the OSEMs do or do not act on the mirror at these frequencies, assuming they don't add extra noise.
4. Page http://nodus.ligo.caltech.edu:8080/40m/5547 seems to indicate OSEM sensor noise is so low it can be neglected.
If we reduce the gain in the OSEMs by adding additional HP filters ( cheby2, HP 0.3Hz, 6dB 4th order ) the happens:
1. MC length gets a bit more noisy at low frequencies - should be looked into some more
2. Coherence between the GUR1 seismometer and MC length goes up between 1E-2 and 1E-1 Hz:
( Ref is with low OSEM gain )
The stacks might be more correlatedly moving together than the pendulums. This would be not so nice for OAF test, but really fine for actually using the MC.
Todo: Measure the OSEM to seismometer coherences with high and low OSEM gains.
For reference the seismometer coherence with one another:
[Jenne, Zach, Frank]
Frank helped Zach and I cable up at PT-100 RTD, and make sure it worked with the Newport Model 6000 Laser Diode Controller. We're using this rather than the Newport 3040 Temperature Controller because Dmass says the output of that isn't working. So we're using just the temp control part of the Laser Diode controller.
The back of the controller has a 15-pin D-sub, with the following useful connections. All others were left Not Connected.
1 & 2 (same) - Pin 2 is one side of TEC output (we have it connected to one side of a resistive heater)
3 & 4 (same) - Pin 4 is the other side of the TEC output (connected to the other side of the resistive heater)
7 - connected to one side of PT-100 temp sensor
8 - connected to other side of PT-100 temp sensor
I used aluminum tape to attach the sensor and heater to the 40m's EOM, and we plugged in the controller. It seems to be kind of working. Zach figured out the GPIB output stuff, so we can talk to it remotely.
The vacuum is ready for no AC power for 1 hr on Sunday morning at 10am
I did the follwing:
Closed V1, stopped the rotation of TP-1 maglev, waited till it reached 0 Hz_ rpm and turned it's controller off.
Closed V4 and stopped TP-2 rotating
Closed all annuloses and VA6
Closed VM1 and opened VM3 This means the RGA is being pumped by TP3. RGA is running in background mode. V5 will close instantly as the AC will be turned off.
VAC STATUS: IFO envelope and annulosses are not pumped. P1 pressure will reach 5-6 mTorr by Sunday morning.
The PSL output shutter will be closed by the interlock at 3 mTorr
Kiwamu will turn off Piezo Jena PZT power supplies and computers Saturday.
I will be here around 1pm Sunday to star pumping. I will need EPICs MEDM running by than.
[Rana, Den, Mirko]
Updated the MC noise projection to include the longitudinal motion of the MC mirrors.
Consistent with static wiener filter showing only benefits in the 1 - 4Hz region.
Jenne gave me a spare LED power switch .
I will replace the broken one on Monday.
By the way here is a picture album of the RF generation box which I took last night.
The power switch button of the RF generation box is not properly working
Another way to make a 1:100 pole:zero boost is to use resistors and capacitors in a Pomona box
mixer -> LB box -> Pomona box -> PZT
Pomona Box = R1 = 7.2 kOhm, C2 = 22 uF, R2 = 72 Ohms (sr560 = $2400, pomona ~ $50)
For the RMS calculation, it would be good to notch out the harmonics. They don't matter since our ALS feedback will have notches at those frequencies.
I wouldn't bother...
I made a super sweet new foam box for our EOM. It's awesome, and should be reasonably easy to duplicate. Check out the PHOTOS!
These are great photos and a nice box, but I fear from the photos that there's too much air getting in. How to pack it so that there's no air flow? How does the temperature sensors wires get in?
We investigated the effect of airflow from the HEPA filters on the PSL beam fluctuation and the resultant noise injected into the WFS loops. The hint that the WFS are injecting PSL beam jitter into MC mirror motion lies in the MC2_TRANS_PIT and YAW signal's power spectrum shown here. First, in the blue trace, which shows the spectrum when the WFS loops are off, we see that the WFS1 and WFS2 error signals have a different shape from that of MC2_TRANS. Since WFS are affected by the PSL beam jitter while the MC2_TRANS_QPD is not, the WFS spectrum contain excess noise, while the MC2_TRANS signals show only the mirror motion. Next, upon switching on the WFS1 and WFS2 loops, we notice that the MC2_TRANS spectra acquire the same shape as the WFS spectra. This shows that the excess noise from the beam jitter has been injected into the MC2 motion, and shows up in the MC2_TRANS spectra.
To confirm these conclusions we repeated the above measurement with the HEPA fans at 0% (Blue trace), 20% (Red), 30% (Brown) and 100% (Green). The plots are shown below. We can see that there is no difference between 0 and 20% levels but beam jitter is visible at 30% HEPA level. The WFS loops were ON during this time and we can can see the PSL noise injected in to MC2 motion (Green).
The HEPA filter fans are now at 20%. How can we be sure that they are really working at 20%, since we cannot see any difference between 0 and 20%?
Now that we have this quiet situation, we also investigated the effect (or lack thereof) of switching on the MC2_TRANS loops. The figure below shows the spectra with all the loops turned off (Blue), with the WFS1 and WFS2 loops turned on (Green) and with everything turned on (Red). With the current output matrix, which is the same simple one as the one in this elog, we see some low frequency suppression. But it also seems to add some noise into the other WFS loops. I am not sure of this result, due the long duration of this measurement, the seimic noise level may have changed over the course of this measurement.
As they are not doing any good just now. I have turned them off by setting the gain in MC2_TRANS PIT and YAW to zero.
[Suresh / Kiwamu]
For tonight we are leaving it as it is but it needs to be fixed at some point.
The temporary solution we decided is to leave it ON so that we can survive tonight.
The box was back in place. The MC is find and 11 MHz and 55 MHz seem okay.
Please be aware of it.
This is a picture showing the rear view of the RF generation box. The red arrow is pointing the blue LED switch button.
Update of the stochmon status
[Attachment 1: Circuit diagram]
- The new stochmon has a low noise amplifier (MAR-6SM) inside.
The RFAM signal from the PD has the power of -60~-50dBm, which is almost at the bottom of the sensitivity for the power detector.
- The band pass filters were doubled.
- I've suffered from some RF coupling from the power line as the power detector is quite sensitive to it.
The situation has been largely improved by the EMI filters in the power supply path, although the problem is still present.
The worst remaining problem is that we can not close the aluminum lid as it cause a huge sprious coupling.
[Attachment 2: Calibration result]
- The outputs were calibarted with Marconi. They showed the signals linear to dBm for the input powers between -70dBm and -10dBm.
- The calibration result was fitted with the empirical fit function. The function and the results are shown in the attachment.
[Attachment 3: Detection limit]
- The attached figure shows the power spectrum of the PD output. This measurement gives us the amount of the RF power given from the PD noise when there is no RF signal.
11MHz out passband noise: −72.7dBm ===> V11 = 2.0483
30MHz out passband noise: −64.6dBm ===> V30 = 1.9333
55MHz out passband noise: −71.2dBm ===> V55 = 2.0272
- Now 11MHz and 55MHz outputs seem indicating the power correctly, but the 29.5MHz output never provides useful information.
It is a constant value independent from the state of the incident beam. Strangely this problem disappears if the marconi is used
for the RF source. Thus this issue is not seen in the calibration measurement.
- So far, 11MHz, 29.5MHz, 55MHz, and DC outputs appear in the channels C1:IOO-RFAMPD_33MHZ, C1:IOO-RFAMPD_133MHZ, C1:IOO-RFAMPD_166MHZ, and C1:IOO-RFAMPD_199MHZ.
They will be renamed.
While the MC was unlocked (and the local damping off) we've measured the coherence between GUR1_X and OSEM sensors. It was rather high, close to 1 at frequencies 0.1 - 1 Hz. That means that stack does not kill all coherence between seismic noise and mirror motion.
Then we've turned on the local damping and measured the coherence again between GUR1_X and OSEM sensors. It decreased due to some noise and was on the level of ~0.5. We did reduced the motion between the mirror and the frame by local damping but it is not obvious that we lost some coherence due to this effect. Probably, actuator adds some noise.
When we locked the MC, we did not see any coherence at 0.1 - 1 Hz between GUR1_X or STS1_X and OSEM sensors of MC1 and MC3 but we did see with MC2. The MC1 sensor was fixed by Suresh.
[Den, Mirko, Suresh]
We were investigating why there is no correlation between MC1 osem signals and seismic motion. During this we noticed a recurrence of this old problem of MC1_LR sensor being dead. I went and pressed down the chip holders where the AA filters used to sit and which now hold the jumper wire. The board is large and flexible it is quite likely some solder joint is broken on the MC1_LR path on this board.
The signal came back to life and is okay now. But it can break off again any time.
Since the MC1 LRSEN channel is not wasn't working, my input matrix diagonalization hasn't worked today wasn't working. So I decided to fix it somehow.
I went to the rack and traced the signal: first at the LEMO monitor on the whitening card, secondly at the 4-pin LEMO cable which goes into the AA chassis.
The signal existed at the input to the AA chassis but not in the screen. So I pressed the jumper wire (used to be AA filter) down for the channel corresponding to the MC1 LRSEN channel.
It now has come back and looks like the other sensors. As you can see from this plot and Joe's entry from a couple weeks ago, this channel has been dead since May 17th.
993190663 = free swinging ringdown restarted again
I modified the Y-Arm PDH box (S/N 17) to have the same TF as the one of the temporary setup described in Kiwamu's earlier entry. Note that the TF below was taken with the gain knob set to 0, so that the proper DC gain is achieved with a setting of ~4. This is desirable because it gives us wiggle room.
The changes were:
Below is the TF along with the LISO model. They are different at low frequencies because the box must have been railing internally (though the phase shows that the result is as expected), and there is a feature around 60 kHz that probably arises from some op amp instability. I will see if adding a small cap somewhere does the trick, and then take a new TF with a lower source voltage.
I'll try to lock the arm with the box tomorrow.
The most interesting plot did not uploaded in the previous elog.
Upload now local MC1_SENSOR signals.
We've found that one of the MC1_SENSORS does not work properly.
See the figure.
MC is unlocked to measure the free swing of the MC mirrors with the local sensors.
Autolocker is disabled.
Analyzing coherence between MC length and signals on MC1, MC2 and MC3 coils we have noticed that MC1 COIL signal is not coherent to MC length at all at interesting frequencies 0.1 - 1 Hz.
We try to explain this phenomena.
8:50PM HEPA@100% for the test
9:20-35PM HEPA level varies from 0%-50%
9:35PM HEPA@40% and left it running at this level
Nov18 1:40 AM HEPA@80% for a work around the PSL table (by KI)
Nov18 4:35 AM HEPA@40% (by KI)
The noise budget on the Y arm ALS has begun.
(Though I feel I made a wrong calibration ... I have to check it again)
The Y arm green PDH servo is working fine with a sufficient amount of suppression.
And the servo configuration looks like this :
(the Error signal)
I took a spectrum of the error signal when the laser was locked to the Y arm and found that it meets the requirement.
Another go at the noise projection from MC1-3 pit/yaw to MC length. This time injecting into the MC autoalignment FB (e.g. C1:IOO-WFS1_PIT_EXC ).
LTPDA is working now, but still the NDS server is not so cooperative.
Summary: Alignment fluctuations of the MC mirrors don't significantly contribute to MC length changes up to at least 3.5Hz. Especially they can't explain the lack of coherence between seismometers and MC length below 1Hz that we worry about for the OAF.
At high frequencies >= 10Hz you can see angle to length coupling as is evident in Sureshes spot position measurements.
Injection from 0.1-20Hz.Filtered by the servo filters and zp:, , Gain = 1 @ 2Hz
Look at the coherence plots for the quality of the measurement:
DOF Amplitude[counts] UTC Time (duration always 4mins)
WFS1p 70 22:28
WFS1y 55 22:03
WFS2p 70 22:13
WFS2y 70 22:18
None - 22:23
To get some better SNR at low frequencies I did a fixed sine noise injection at 0.3Hz. See attached files.
DOF Amplitude[counts] UTC Time (duration always 4mins) Lower limit of SNR MC length via mirror misalignment
WFS1p 4 00:05 29.3
WFS1y 4 00:14 22.0
WFS2p 4 00:19 18.5
WFS2y 4 00:25 18.0
Our existing 300 series SS plungers from McMastercar #8476A43 are silver plated as Atm2 shows.
Problems: 1, they become magnetized after years being close to the magnets
2, they oxidize by time so it is hard to turn them
I looked around to replace them.
Titanium body, nose and beryllium copper spring. None magnetic for UHV enviorment.
Can be made in 7 weeks at an UNREASONABLE $169.00 ea at quantity of 50
It turns out that we don't have all the parts I would need to do a full prototype of the precision temperature controller. I am guessing that we won't want to sit around and wait for the parts given the upcoming TAC meeting, so I'll do the next best thing:
Does anyone have a suggestion for how this thing will be packaged? I.e., should it be in a box or should it be mounted in a rack, etc. In the end, a real board will be printed and stuffed, so this need not be a really professional job in the short term.
Is there an update on Stochmon? Are the signals acquired somewhere already? What's the current deal-io? The new EOM mount should be here later today, and I'm jazzed to start checking how my EOM box helps (hopefully) the amount of RFAM we see.
I'll start making the adapter plate while I wait...
Actually, do we need to reset the filter history at every lock loss of the MC?
Those DC offsets were necessary to keep the alignment good just until the MC is unlocked.
So if we keep the history, we can maintain the good alignment.
I suspect the integrators get fed a huge wrong signal on lockloss. Clearing the history on the trans DOFs when the MC was badly aligned gets it nicely aligned again. I switched off the alignment transmission DOFs for now.
I have modified the 'mcwfson' and 'mcwfsoff' scripts to include the Clear History step for the MC2_TRANS_PIT and _YAW filters.
These scripts can be run, by hand, from LOCKMC screen or from the WFS_MASTER screen. Use the 'Turn WFS ON/OFF' button.
The mcautolockmain script will now clear history on all ASC filter banks when the MC unlocks.
I have turned on ASC loops on the MC2_TRANS (= alignment transmission DOFs of the above elog) paths.
I'm just on an elog roll this morning...
Again while poking around inside the IFO room, I noticed that they have replaced all of our incandescent lights with CFLs. Do we care? The point of having the incandescent lights on a separate switch from the big fluorescent lights was so that we could have only 60Hz lines, not wide-band noise if we want the lights on while locking.
I'm not sure that we actually care, because more often we just turn off all the lights while trying to do serious locking, but if we do care, then someone needs to ask the custodial staff (or someone else?) to undo the change.